4. A racer accelerates from rest at a constant rate of 2 m/s2. (a) How fast will the racer be going at the end of 6.0 s? (b) How far has the racer traveled during this time?




(Answer: a. 12 m/s & b. 36 m)

1
Expert's answer
2022-05-30T15:17:38-0400

(a)

"v=v_0+at\\
=0+2*6=12\:\rm m/s"

(b)

"d=v_0t+at^2/2\\
=0*6+2*6^2/2=36\:\rm m"